OpenStudy (anonymous):

kelly has a total of 30000 to be invested in two municipal bonds. Bond A yields 8% interest per year and bond b yields 10% interest per year. How much should kelly invest in each bond if she requires an annual income of 2740 from her bond investments?

OpenStudy (anonymous):

x+y=30000 8%x+10%y=2740 x+y=30000 y=30000-x 8%x+10%y=2740 0.08x+0.1y=2740 0.08x+0.1(30000-x)=2740 0.08x+3000-0.1x=2740 -0.02x=2740-3000 -0.02x=-260 x=13000(this is an amount that has 8% interest) x+y=30000 13000+y=30000 y=30000-13000 y=17000(this is for an investment with 10% interest)

OpenStudy (anonymous):

Thank you!!